Complete the ordered pairs so that each is a solution for the given problem: 3x+4y=12 (_,0),(8/3,_)
Answer by funmath(2873)

Complete the ordered pairs so that each is a solution for the given problem: 3x+4y=12 (_,0),(8/3,_)
Ordered pairs are in alphabetical order (x,y)
For (_,0) let y=0 and solve for x.
3x+4(0)=12
3x+0=12
3x=12
3x/3=12/3
x=4 So the ordered pair is (4,0)
For (8/3,_) let x=8/3 and solve for y
3(8/3)+4y=12
8+4y=12
8-8+4y=12-8
4y=4
4y/4=4/4
y=1 So the ordered pair is (8/3,1)
Happy Calculating!!!!